OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope -2/3 and y-intercept (0, 3) and graph the line. please show work and show graph!

OpenStudy (anonymous):

whats y intercept?

OpenStudy (anonymous):

(0,3)

OpenStudy (dumbcow):

y-3 = -2/3(x-0) y = -2/3x +3

OpenStudy (dumbcow):

to graph start at (0,3) then go down 2, over 3

OpenStudy (anonymous):

thanks so much can u help me with a couple of more problems!

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

hold on give me three mins to post them ok!